1. A scientist wants to store 54.00 kg of krypton gas, K(g) in a 80.00 L container in a refrigerated  4.00°C room. What pressure capacity, in psi, will the container need to have? {/6}
1
Expert's answer
2021-02-09T03:37:45-0500

According to the formula PV=nRT:

P=nRT/V.

n=m/M=54000g/84g*mol-1=642.86mol.

For volume in litres R=0.082.

T=4.00°C+273=277.00K.

P=642.86*0.082*277.00/80.00=182.52(atm).

P=182.52*14.70=2683.04(psi)